answersLogoWhite

0

The probability of drawing a heart in a single random draw from a normal deck of playing cards and rolling a 4 on a fair number cube is:

1/4*1/6 = 1/24

User Avatar

Wiki User

∙ 12y ago

What else can I help you with?

Related Questions

What is the probability of drawing a heart or club from a pack of cards?

Probability of drawing a heart: 1/4 Probability of drawing a club: 1/4 Probability of drawing a heart or a club: 1/4 + 1/4 = 2/4 = 1/2


What is the probability of getting a red heart in a deck?

The probability of drawing a red heart is 1 in 4. This is the same as the probability of drawing a heart, as red is included as a superset of hearts.


What is the probaaility of drawing a face card that is not a heart?

There are 12 face cards and three of them are hearts. Therefore, the probability of drawing a face card that is heart is 3/12 = 1/4. The probability of drawing a face card that is not a heart is 1-1/4 = 3/4.


What is the probability of drawing two hearts from a standard pack of cards is 3 divided 51 what is the probability that the two cards drawn are not both hearts?

The probability of drawing a heart from a fair deck is 1 in 4. If the card is replaced then the probability is again 1 in 4. The probability of drawing a card other than a heart is 3 in 4. Once again if the card is replaced then the probability remains 3 in 4


Drawing a heart from a deck of cards?

There are only 4 suits in an ordinary deck of playing cards. The probability of drawing a heart is 1 in 4, or 0.25.


A single card is drawn from a deck Find the probability of selecting a heart or a 9?

The probability of drawing a heart is 13 in 52 or 1 in 4. The probability of drawing a 9 is 4 in 52 or 1 in 13. The probability of drawing a heart or a 9 is (13 + 4 - 1) in 52, with the -1 required because the 9 of hearts intersects the solution set. (13 + 4 - 1) in 52 is 16 in 52, which is 4 in 13, or about 0.3077.


What is the probability of not drawing a heart?

Number of cards in a deck = 52 Number of cards that are heart = 13 Therefore number of cards that are not heart = 52-13 = 39 Probability of not drawing a heart = 39/52 or 3/4


What is the probability of drawing either a 3 or a heart from a regular bridge deck of cards?

The probability of A is denoted P(A) and the probability of B is denoted P(B). P(A or B) = P(A) + P(B) - P(A and B). Say P(A) = Probability of drawing a heart, which is 13/52. Say P(B) = Probability of drawing a three, which is 4/52. We now have to determine P(A and B) which is the probability of a heart and a three, which is 1/52. We now can determine the probability of drawing a heart or a three which is 13/52 + 4/52 - 1/52 = 16/52 = 4/13.


What is the the probability of rolling a 4?

The probability of rolling a 4 is 1/6.


What is the probability of pulling a heart out of a deck of cards?

The probability of drawing a heart in a standard deck of 52 cards is 13 in 52 or 1 in 4.


What is the probability of drawing a heart?

! in 4, as the four suits have an equal number of cards.


If one card is drawn from a standard deck of 52 playing cards what is the probability of drawing a heart?

There are 13 hearts in a 52 card deck so the probability of drawing a heart is 13/52 or 1/4.

Trending Questions
Where is the deck of cards in where's waldo deep-sea divers? What percentage of men are impotent all their life? A bowl contains a total of 26 blue and red marbles There are two more red marbles than blue marbles How many blue marbles are in the bowl? What is the difference between dart rate and severity rate? In a report to show a correlation between a full moon and animal bites which table would be most appropriate? How do you find the sample size if you are given the confidence interval and the margin of error as well as the standard deviation? An ordinary six-sided dice is thrown what is the probability that the score is greater than 3? Are tornadoes fairly common in the state of Texas? If he pulls out after sex how likely are you to get pregnant? What is the differences between subjective risk and objective risk? Is the square footage of a house discrete or continuous? The decimal form 1.46 becomes? What is an example of quantitative and qualitative data? A drawer cotains 50 bolts and 150 nuts. Half the of the bolts are rusted if one item is chosen at random what is the probability that it is rusted or a bolt? What is the factor in premium computation that has to do with record keeping and statistical analysis insurance companies perform? What difference between data warehouse and traditional database used OLTP? Is the number of cars owned discrete or continuous? What is the average weight for a 6' 1'' 19 year old? What is the perimeter of a room? How do you Find the mean with a random variable and a probability?